# Beyond Conventional Transformers: The Medical X-ray Attention (MXA) Block for Improved Multi-Label Diagnosis Using Knowledge Distillation

### v1

@@ Abstract
Keywords Computer Vision  $\cdot$ Medical Imaging  $\cdot$ Multi-Label Classification  $\cdot$ Vision Transformers  $\cdot$ X-ray Diagnosis
@@ 1 Introduction
The adaptation of the attention-based transformer architecture [1] for computer vision, leading to the development of Vision Transformers (VITs) [2], demonstrates exceptional performance in tasks such as image classification [2], segmentation [3], and generation [4]. The ViT showcases superior model performance in contrast to traditional methodologies such as Recurrent and Convolutional Neural Networks [2]. However, these advances in accuracy came with the downside of an increased model size and computational cost.
Important breakthroughs have been made in medical analysis with the adoption of deep learning techniques, particularly CNNs [5] and transformers. CNN-based architectures like U-Net [6] have set benchmarks in segmentation tasks, enabling precise delineation of organs and anomalies in medical images. However, non-CNN architectures such as transformers and hybrid models (e.g., TransUNet [7]) have recently gained traction, leveraging the global attention capabilities of transformers to achieve superior performance in tasks like segmentation. For instance, Swin Transformers [8], with their hierarchical structure and shifted windowing mechanism, have been instrumental in improving the efficiency and scalability of vision models, making them suitable for tasks requiring high-resolution input. Additionally, models like CheXNet [9], trained on chest X-rays, have been pivotal in diagnosing diseases like pneumonia with performance comparable to radiologists.
Despite the advancements in transformer-based architectures for vision tasks, their practical deployment in clinical domains like medical imaging faces significant challenges. Transformers typically require extensive computational resources and large-scale datasets for training, both of which are often scarce in medical imaging [10, 7]. This limitation is particularly pronounced in X-ray analysis, where datasets are typically smaller, and many publicly available datasets suffer from issues such as inconsistent data collection procedures and biased performance assessment metrics, leading to poor generalization on unseen data [11, 12, 13]. These challenges underscore the need for models that are both computationally efficient and capable of handling the inherent data limitations in medical imaging.
Efforts to address these challenges have led to the development of efficient transformer architectures, such as Data-efficient Image Transformers (DeiTs) [14] and EfficientViTs (EfficientViTs) [15], which reduce computational overhead while maintaining competitive performance. For instance, DeiTs leverage knowledge distillation to improve efficiency, while EfficientViTs optimize the balance between Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN) layers to achieve significant gains in throughput and accuracy. These advancements make them particularly suited for medical imaging applications, including X-ray analysis, where computational resources are limited and scalability is critical. The shift from CNNs to pure transformer architectures is driven by the state-of-the-art performance of Vision Transformers (ViTs) in vision tasks such as ImageNet classification [2], highlighting their immense potential contribution to the field of medical imaging.
Additionally, multi-label classification is a crucial yet underexplored area in X-ray diagnostics, as a single scan often reveals multiple conditions simultaneously [9, 16]. In this work, we propose a framework to address these gaps. Specifically, our contributions are as follows:
  1. 1.
    Multi-Label EfficientViTs: We adapt the EfficientViT (EfficientViT) model [15] to handle multi-label outputs. By replacing cross-entropy loss with BCEWithLogitsLoss and introducing CheXpert [16], our approach efficiently handles multiple pathologies within a single forward pass.
  2. 2.
    Medical X-ray Attention (MXA) Block: We introduce the MXA block, which integrates Dynamic Region of Interest (ROI) Pooling [17] and CBAM [18]-style attention in parallel with the standard Multi-Head Self-Attention layers. This aims to capture both localized abnormal findings (often subtle in chest X-rays) and global context without incurring large computational overhead.
  3. 3.
    Multi-label Knowledge Distillation: To further boost performance under data constraints, we employ a multi-label distillation scheme. A teacher model (e.g., DenseNet-121 [14] pre-trained on CheXpert [16]) provides soft or hard labels, thereby improving the student transformer’s recognition of difficult pathologies.
We validate our method on CheXpert, demonstrating that our multi-label EfficientViTs with the MXA block achieve improved accuracy and computational efficiency. Notably, the parallel integration of MHSA and MXA—augmented with knowledge distillation [19]—enables both robust global awareness and fine-grained local feature extraction, addressing common pitfalls in chest X-ray interpretation.
The rest of the paper is organized as follows: we review related works in Section 2, including advancements in multi-label image classification and medical imaging techniques. Section 3 details our approach to adapting EfficientViTs for multi-label classification. Section 4 introduces the Medical X-ray and Section 5 outlines our approach to Knowledge Distillation. Section 6 explains our experimental setup, including the dataset, evaluation metrics, and implementation details. Section 7 presents the results and the implications of our findings, limitations, and future directions for research in transformer-based medical imaging models.
2 Related Work
2.1 Multi-label Image Classification
Image classification has been a key application of computer vision, with substantial advancements driven by deep learning. Convolutional Neural Networks (CNNs) [5] have historically been the backbone of image classification, achieving state-of-the-art performance on datasets like ImageNet [20]. Architectures such as AlexNet [21], ResNet [22], and EfficientNet [10] have progressively improved model performance by introducing innovations in depth, width, and computational efficiency. More recently, Vision Transformers (ViTs) [2] have demonstrated the ability to outperform CNNs by leveraging self-attention mechanisms, particularly in large-scale datasets, thereby highlighting the potential of transformer-based architectures in image classification.
While significant progress has been made in single-label image classification, multi-label image classification presents additional challenges. Unlike single-label tasks, where each image is associated with a single category, multi-label classification requires models to predict multiple, potentially overlapping labels for a single image. This is particularly important in real-world applications such as clinical settings of medical imaging, where a single scan can reveal multiple conditions. Standard deep learning architectures, such as ResNet and DenseNet, have been adapted to multi-label classification by employing sigmoid non-linearity functions and binary cross-entropy loss to handle multiple labels independently [23, 24].
In the medical imaging domain, multi-label classification has been explored for tasks such as diagnosing chest X-rays, where an image may indicate multiple abnormalities. CheXNet [9] was among the first models to utilize a deep learning-based approach for chest X-ray classification, achieving radiologist-level performance on pneumonia detection while identifying other conditions. Later works, such as those leveraging the CheXpert dataset [16], highlighted the challenges of handling label uncertainty and imbalance in multi-label datasets. Transformers have recently been explored in this context due to their ability to model global dependencies, making them well-suited for identifying co-occurring conditions. For example, TransMed [25] applies transformer-based models for multi-label classification in the medical domain, demonstrating their potential in improving diagnostic accuracy.
2.2 CheXNet: Pneumonia Detection in Chest X-rays
CheXNet [9] is a pioneering work in the application of deep learning to chest X-ray analysis. This model utilizes a 121-layer DenseNet [26] architecture to perform multi-label classification, diagnosing up to 14 thoracic diseases. The dataset used for training was the NIH Chest X-ray dataset [27], which contains more than 100,000 images labeled with multiple conditions. The primary innovation of CheXNet is its ability to achieve radiologist-level performance on pneumonia detection, demonstrating the efficacy of CNNs [5] for X-ray diagnostics.
Despite its success, CheXNet is constrained by its reliance on large-labeled datasets, and its performance can degrade when faced with noisy or unbalanced data. Furthermore, while CheXNet addresses multi-label classification, it does not explicitly model dependencies between labels, which limits its ability to handle co-occurring conditions effectively.
Building on these limitations, we aim to leverage the power of transformer-based architectures to further improve performance in X-ray diagnostics. By combining methodologies such as knowledge distillation and cascading group attention, we aim to achieve higher accuracy in multi-label classification while introducing methods to reduce computational overhead.
2.3 CheXpert: A Benchmark for Chest X-ray Diagnosis
CheXpert [16] is a large-scale dataset designed to benchmark the performance of machine learning models in chest X-ray diagnosis. It consists of over 224,000 labeled chest radiographs from 65,000 patients, annotated with uncertainty labels for 14 common thoracic conditions. CheXpert addresses some of the challenges faced by prior datasets, such as label imbalance and inconsistency, by employing a rule-based labeling system to extract observations from radiology reports. This enables the creation of a standardized dataset that serves as a robust benchmark for evaluating machine learning models.
The CheXpert dataset has driven innovation in chext X-ray analysis. Current top-performing approaches often integrate advanced architectures, such as DenseNets and transformers, with tailored loss functions and training strategies. For example, in "Large-scale Robust Deep AUC Maximization" [28], a novel surrogate loss function is introduced designed to maximize the Area Under the Curve (AUC) for imbalanced datasets. This method, tested on CheXpert, achieves state-of-the-art results by directly optimizing AUC, a metric critical in medical imaging tasks.
Despite these advances, the computational cost associated with these models remains a challenge, particularly for real-time or resource-constrained applications. Our work builds upon these advancements by introducing EfficientViTs coupled with knowledge distillation, aiming to improve computational efficiency while maintaining state-of-the-art performance on benchmarks such as CheXpert.
2.4 DenseNet-121
The Dense Convolutional Network (DenseNet) architecture was proposed by Huang et al. [26] to improve information flow in convolutional neural networks by introducing dense connectivity patterns. Unlike traditional CNNs, where each layer only connects to the following layer, DenseNets connect each layer to every other layer in a feed-forward manner. This design drastically enhances gradient propagation and feature reuse throughout the network, enabling models to be both deeper and more parameter-efficient.
DenseNet-121 [26] specifically has 121 layers organized in dense blocks separated by transition layers, striking a good balance between accuracy and computational cost. In medical imaging, DenseNet-121 has been especially popular due to its ability to learn rich feature representations from relatively limited datasets. For instance, Rajpurkar et al. [9] employed a DenseNet-121 backbone in CheXNet—one of the pioneering works that applied deep learning to chest X-rays—achieving near-radiologist-level performance in detecting pneumonia. Subsequent adaptations of DenseNet-121, often fine-tuned on large chest radiograph datasets (e.g., NIH ChestX-ray14 [27] or CheXpert [16]), have shown state-of-the-art results for a variety of thoracic abnormalities.
3 Multi-labeling with Efficient Vision Transformers
While the original EfficientViTs [15] implementation was geared towards single-label tasks (e.g. ImageNet classification), we extend its functionality to multi-label classification, an essential capability in clinical X-ray diagnosis, where multiple abnormalities may coexist. This section details our modifications to EfficientViTs, including the transition from cross-entropy (CE) to multi-label loss, changes to the dataset loader, and the integration of a multi-label knowledge distillation configuration.
3.1 Adjusting the Loss Function and Output Layer
In single-label classification, the training loop typically relies on a CE loss [29] with softmax outputs. However, multi-label classification requires each sample to predict a set of labels independently. Therefore, we replace the single-label objective with Binary Cross Entropy with Logits (BCEWithLogitsLoss), which effectively combines a sigmoid activation and binary cross-entropy in one step.
Let $o=(o_{1},...,o_{C})$ be the raw logits output by the network for $C$ classes, and let $y=(y_{1},...,y_{C})$ be the multi-label ground truth, where each $y_{C}\in[0,1]$. BCEWithLogitsLoss is defined as:
$\mathscr{L}_{\text{BCEWithLogits}}(o,y)=-\sum_{c=1}^{C}[y_{c}\text{log}(\sigma(o_{c}))+(1-y_{c})\text{log}(1-\sigma(o_{c}))]$
where $\sigma$ is the sigmoid function. In our case, we set $C=14$ to match the CheXpert dataset [16]. After the final pooling layer, we thus produce logits of shape $(B,14)$ and if multi-labeling is active, we apply $\sigma$ to map logits to probabilities $[0,1]$ for each pathology.
3.2 Dataset and Data Loading for Multi-label Targets
In contrast to single-label datasets (e.g., ImageNet), our CheXpert dataset [16] provides annotations for up to 14 possible pathologies per image. We implement a custom CheXpertDataset class that reads the CheXpert CSV and constructs a multi-label vector for each X-ray. Specifically, for each row we parse 14 binary indicators (e.g., Atelectasis, Cardiomegaly, Pneumonia) and store them in a tensor of shape (14,). We follow the “U-1” protocol, treating uncertain (U) labels as positive (i.e., 1) rather than negative (0). This choice is motivated by findings in [16] and confirmed by Table 2 in [30], which show improved performance when uncertain labels are approximated as positive. Formally, for each pathology label:
$y_{c}=\begin{cases}1,&\text{if the label is 1 or uncertain},\\ 0,&\text{otherwise}.\end{cases}$
Though uncertain annotations can be mapped to 0, 1, or even ignored entirely, recent work suggests that “U-1” often yields better diagnostic performance than “U-0,” especially for pathologies where missing or subtle features might still be clinically relevant [30, 11]. Our approach is straightforward yet effective, and can be adapted depending on domain-specific requirements.
This design integrates well with the multi-label BCEWithLogitsLoss, allowing the network to learn from both definitively positive and uncertain/ambiguous labels in parallel. As discussed further in Section 3.3, we feed these labeled samples into the EfficientViT training loop, ensuring that each of the 14 possible pathologies can be predicted independently per image.
3.3 Training and Evaluation
Our training loop (adapted from EfficientViTs) still leverages the original DistillationLoss wrapper but replaces the base cross-entropy criterion with BCEWithLogitsLoss. Let $B$ be the batch size and let $O\in\mathbb{R}^{B\times 14}$ be the output logits. The multi-label ground truth is $Y\in[0,1]^{B\times 14}$. During each forward pass,

### v2

@@ 1 Introduction
Chest X-rays (CXR) are the most common medical images, yet one scan often reveals several co-existing pathologies that automated readers must detect simultaneously. Vision Transformers (ViTs) [1] excel at global reasoning, but their quadratic self-attention makes deployment on high-resolution CXR costly in both parameters and floating-point operations (FLOPs). Pure convolutional networks, e.g. DenseNet-121 [2], run fast but miss long-range cues spread across the thorax. The clinical setting therefore demands models that (i) learn multi-label dependencies, (ii) highlight subtle local findings, and (iii) fit the limited compute of hospital workstations.
Recent efficient transformers, DeiT [3], Swin [4], EfficientViT [5], trim FLOPs via knowledge distillation (KD) or windowed attention, yet none are specialized for medical images where abnormalities occupy just a few pixels. Moreover, public CXR datasets such as CheXpert [6] remain modest by ViT standards, amplifying over-fitting and class imbalance. We address these gaps with a domain-aware transformer that pairs global self-attention with inexpensive region attention and learns from a calibrated CNN teacher.
Contributions.
Our work makes three principal contributions: (i) We extend the EfficientViT to multi-label chest-X-ray diagnosis by replacing the cross-entropy objective with a class-balanced BCE with logit loss, enabling a single forward pass to predict all 14 CheXpert labels. (ii) We introduce the lightweight Medical X-ray Attention (MXA) block, which pairs dynamic ROI pooling [7] with channel–spatial gating [8] in parallel to multi-head self-attention, capturing subtle lesions without increasing asymptotic complexity. (iii) We devise a multi-label knowledge-distillation scheme in which a frozen DenseNet-121 teacher provides soft targets, guiding the transformer on rare pathologies and stabilizing training on the 224 k-image CheXpert corpus [9].
On CheXpert our EfficientViT-MultiLabel-M5 model reaches 0.85 mean ROC-AUC, improving the M5 EfficientViT baseline by +0.19 while preserving its 5.7G FLOPs footprint. MXA alone delivers half the gain; distillation adds the remainder, confirming their complementary roles. The resulting architecture offers an attractive accuracy-to-efficiency trade-off for point-of-care deployment and illustrates how task-specific attention bridges the gap between generic ViTs and radiologist needs.
2 Related Works
Multi-label image classification and CheXNet.
Deep CNNs, from AlexNet to EfficientNet, drove single-label ImageNet accuracy upward, but Vision Transformers (ViTs) surpassed them once large-scale data or distillation became available [10, 11, 12, 1, 3]. In multi-label chest X-ray diagnosis several findings (e.g., consolidation, edema, cardiomegaly) often co-occur. The per-label sigmoid activation followed by Binary Cross-Entropy (BCE) loss, widely adopted for such tasks, optimizes each label in isolation. This independence assumption disregards inter-label structure and tends to underweight infrequent diseases, leading to poorer AUC on rare classes [13, 14]. CheXNet mitigated class imbalance by fine-tuning DenseNet-121 on 112 k NIH scans and achieved radiologist-level pneumonia detection [2]; however, its purely convolutional backbone still treated pathologies independently and lost accuracy on external cohorts. Transformer variants such as TransMed explicitly model label dependencies, yet their quadratic self-attention inflates FLOPs, limiting bedside deployment [15]. Our approach grafts a lightweight Medical X-ray Attention block onto EfficientViT [5] and distills knowledge from a calibrated DenseNet teacher, preserving CheXNet’s clinical strengths while delivering ViT-level global reasoning at workstation-level cost.
CheXpert benchmark.
CheXpert offers 224k radiographs from 65k patients with rule-mined, uncertainty-annotated labels for the same 14 findings, providing a stronger test bed than NIH for real-world deployment [6]. Its public leaderboard has steered advances ranging from label-uncertainty strategies (e.g., U-on-1, self-training) to loss designs that directly maximize AUC on imbalanced data [16]. Transformer backbones, including DeiT and Swin, now dominate the leaderboard but often rely on extensive augmentations and multi-crop inference that inflate latency. However, computationally heavy models impede clinical application where GPUs may not be available. We thus adopt CheXpert as our primary benchmark yet target an efficiency regime suitable for clinical integration: EfficientViT halves FLOPs versus Swin while our MXA block and distillation recover the accuracy gap. By evaluating under identical U-1 labeling and augmentation settings, we isolate the benefit of task-specific attention rather than dataset heuristics.
DenseNet-121 as teacher.
DenseNet-121’s dense skip connections deliver strong feature reuse with only 8M parameters [17]; consequently it is the default backbone in many open-source chest-X-ray libraries such as TorchXRayVision. When pretrained on CheXpert the model attains robust AUC across diverse pathologies and exhibits calibrated probability outputs, making it an ideal teacher for KD. Prior student–teacher pairs have compressed DenseNet into MobileNet or ShuffleNet for on-device screening, but rarely into transformers [18, 19]. We keep the teacher frozen and distill its logits into an EfficientViT student, showing that soft multi-label distillation narrows the performance gap while letting the transformer exploit global context unavailable to the convolutional teacher.
3 Methods
3.1 Multi-labeling with Efficient Vision Transformers
While Cream’s EfficientViT [^1] https://github.com/microsoft/Cream/tree/main/EfficientViT was engineered for single-label tasks (e.g. ImageNet classification) [5], CXR diagnosis demands multi-label predictions, as several abnormalities can co-occur. To handle the co-occurrence of multiple pathologies in a single chest X-ray, we adapt EfficientViT’s single-label framework by redesigning its output head and training loss for multi-label prediction. Specifically, we replace the softmax head with a per-class linear layer trained under a class-balanced binary cross-entropy objective, enabling one efficient transformer pass to simultaneously predict all target abnormalities.
Label representation & loss objective.
We follow the “U-1” protocol (Appendix A), motivated by best practice  [6] [20]. EfficientViT’s final pooling layer is replaced by a $(B,C)$ linear head whose logits $o=(o_{1},\dots,o_{C})$ are trained with Binary Cross-Entropy with Logits (BCEWL); full definition in Appendix B. The formulation naturally accommodates independent per-pathology predictions, while letting the network exploit both definite and uncertain positives, and yields probability scores $\sigma(o_{c})\!\in\![0,1]$ used for AUC and threshold accuracy metrics.
Training, Loss & Evaluation.
We train our EfficientViT student on batches of size $B$ by minimizing a convex combination of the ground-truth multi-label BCEWL and a soft-label knowledge-distillation term from our frozen DenseNet-121 teacher. Specifically, with student logits $O^{s}\in\mathbb{R}^{B\times 14}$, true labels $Y\in[0,1]^{B\times 14}$ under the U-1 protocol (Appendix A), and teacher probabilities $p^{t}=\sigma(O^{t})$, our total loss is
$\mathscr{L}_{\text{total}}\;=\;(1-\alpha)\,\mathscr{L}_{\text{BCEWL}}\!\bigl{(}O^{s},Y\bigr{)}\;+\;\alpha\,\mathscr{L}^{\text{soft}}_{\text{KD}}\!\bigl{(}O^{s},p^{t}\bigr{)},$
where $\alpha\!\in\![0,1]$ controls the distillation strength. During each forward pass we compute $O^{s}=\mathrm{EfficientViT}(\text{samples})$ and back-propagate through the student only. At validation time we evaluate the exact same loss and additionally measure two key metrics: (1) mean area under the ROC curve (AUC) across all 14 pathologies, which captures discrimination across thresholds, and (2) per-label accuracy at a 0.5 probability cutoff, as a proxy for binary clinical decision performance, by thresholding $\sigma(O^{s})$ and comparing to $Y$. All reported metrics are averaged over three random seeds, with 95% confidence intervals where applicable. Our primary objective is to boost mean AUC over the vanilla EfficientViT baseline, especially on under-represented diseases, while maintaining or exceeding baseline threshold accuracy.
3.2 Medical X-ray Attention (MXA) Block
Figure 1: Integration of the Medical X-ray Attention (MXA) Block Architecture. The MXA block injects ROI pooling + CBAM gating in parallel with MHSA, focusing compute on abnormal regions.
Dynamic region-of-interest (ROI) pooling.
Multi-label X-ray classification often includes multiple abnormal regions and purely global feature pooling might dilute small or subtle findings. To identify the most relevant regions for targeted analysis, consequently reducing computational overhead, we propose a Dynamic ROI pooling approach. By dynamically pooling region-of-interest (ROI) features, the model sub-sectionalizes critical areas while ignoring extraneous regions.
We employ a learnable ROI-pooling module in which a lightweight convolutional predictor generates normalized bounding-box coordinates $\mathrm{ROI}_{i}=[x_{1},y_{1},x_{2},y_{2}]$ for each feature map $F\in\mathbb{R}^{B\times C\times H\times W}$. These boxes are used to crop and rescale $F$, and the multi-label classification loss back-propagates through the ROI predictor so that box localization and classification parameters are optimized end-to-end (see Appendix C for full details). After cropping the predicted ROI from the feature map, we resize it to the original feature-map size via bilinear interpolation:
$F^{(i)}_{\text{pooled}}\;=\;\operatorname{Resize}\!\bigl{(}F_{\mathrm{ROI}_{i}},\,(H,W)\bigr{)},$
where $F^{(i)}_{\text{pooled}}$ is the pooled ROI feature map and $\operatorname{Resize}(\cdot)$ denotes bilinear interpolation. By pooling only the relevant region to full resolution, we reduce the overhead that would otherwise come from processing the entire X-ray at high granularity [21]. This ensures compatibility with subsequent operations while focusing on the most critical regions. We use a kernel size of 3 in certain layers (e.g. local window attention or ROI-based convolutional components) as a practical middle ground for ROI between capturing sufficient local context and avoiding over-fitting in CXR data. Prior work in medical image analysis (e.g., [22]) has similarly found that intermediate kernel sizes strike a good balance between detail preservation and robust feature extraction.
Convolutional Block Attention Module (CBAM)-style attention.
To further refine the representation and enhance localized feature learning, we incorporate a CBAM [8] into the MXA block. CBAM sequentially applies channel attention and spatial attention to the pooled feature maps, enabling the network to focus on the most informative channels and spatial regions. This combined attention helps the network prioritize more pathologically relevant features, crucial in medical images where anomalies can be small or faint.
Channel attention determines which feature channels contribute the most to the classification task. This is achieved by globally aggregating the feature map along the spatial dimensions using global average pooling (GAP) and global max pooling (GMP), followed by a shared multi-layer perceptron (MLP) [23]; see Appendix D for the formal definition.
The combined effect of channel and spatial attention refines the feature representation by prioritizing both the most informative channels and regions. Incorporating CBAM into the MXA block thus provides a richer and more adaptive representation of X-ray feature maps.
Parallel integration with MHSA.
To further enhance feature representation and ensure that the model effectively utilizes both global dependencies (captured by MHSA [24]) and localized features (captured by MXA), we propose integrating the MXA block in parallel with the MHSA layers in EfficientViTs.
$F_{\text{output}}\;=\;F_{\text{MHSA}}\;+\;F_{\text{MXA}},$
where $F_{\text{MHSA}}$ represents the output from the MHSA layer, and $F_{\text{MXA}}$ is the refined feature map from the MXA block. Refer to Appendix F, Fig. F.1 for an illustration of the MXA block and its integration into the overall transformer framework.
3.3 Knowledge Distillation
Transformation layer and label mapping.
Knowledge distillation (KD) transfers knowledge from a large, well-trained teacher to a compact student model [25]. Within our EfficientViT framework for multi-label CXR classification, we use the TorchXRayVision [^2] https://github.com/mlmed/torchxrayvision DenseNet-121 teacher (specifically the “densenet121-res224-chex” variant) [17, 2, 26], pretrained on CheXpert [6] and producing 18 logits. Our student predicts 14 pathologies, so we insert a lightweight adapter to align the teacher’s 18-dimensional output with the 14 labels expected by the student. Let $O^{\text{t}}\in\mathbb{R}^{18}$ denote the teacher logits. First, a fixed index map $M$ permutes the logits so that each student index $k$ receives the correct teacher value (e.g., $M(1)=17$). Second, because the student includes a No Finding class that is absent in the teacher, we synthesize its probability by combining the teacher’s per-pathology scores:
$p_{\textsc{NF}}=\prod_{i=1}^{18}\bigl{(}1-\sigma(O^{\text{t}}_{i})\bigr{)},\qquad\text{logit}_{\textsc{NF}}=\sigma^{-1}(p_{\textsc{NF}}).$
This quantity captures the likelihood that no abnormality is present. Finally, logits corresponding to teacher labels with no student counterpart (e.g., Support Devices, Pleural Other) are set to zero so they receive no gradient. This procedure reconciles label-space mismatches and passes a clean 14-dimensional target to the student, allowing KD to sharpen recognition of subtle or co-occurring findings.
Soft distillation and dynamic weighting.
The adapted teacher probabilities from the original $O^{\text{t}}\in\mathbb{R}^{18}$ logits are represented as $p^{\mathrm{t}}=\sigma(O^{\mathrm{t}})$. We adopt soft distillation because it preserves the teacher’s uncertainty. To focus learning where the teacher is least confident, we introduce dynamic label weights